OpenStudy (anonymous):

Find the integral t(t^2+7)^2

OpenStudy (amistre64):

you could expand it out; or just notice that your missing a 3 and a 2

OpenStudy (anonymous):

I am confused by what you just said the question is \[\int\limits_{}t(t^2+7)^2\]

OpenStudy (anonymous):

with dt after it

OpenStudy (amistre64):

well, if we notice this as a x^2 type format; then we should see that it comes from something that looks like x^3/3 lets try that out and see what we get when we derive it (t^2+7)^3 --------- derives to what? 3

OpenStudy (anonymous):

t^2+7

OpenStudy (amistre64):

not quite; bring down the exponent, subtract the 1 and pop out the derivative of the innards

OpenStudy (anonymous):

(t^2+7)^2

OpenStudy (amistre64):

good, and the chain rule pops out a derivative of the inside parts

OpenStudy (amistre64):

otherwise: t(t^2+7)^2 = t (t^4+14t^2+49) = t^5 +14t^3 + 49t \[\int t^5 +14t^3 + 49t\ dt\] is just as usable

OpenStudy (amistre64):

most likely this is an example of a u-substitution .... which isnt needed but might help to claen things up

OpenStudy (anonymous):

ya this chapter is over the substitution

OpenStudy (amistre64):

say u = t^2 + 7 du/dt = 2t du/2 = t dt \[\int \frac{u^2}{2}du\]

OpenStudy (amistre64):

in this format, its just a simple integration that you should be used to by now

OpenStudy (anonymous):

This is the step I get lost at

OpenStudy (amistre64):

the u subbing or the integration? :)

OpenStudy (amistre64):

its a method that changes the way the integrand looks, while keeping the same value

OpenStudy (amistre64):

replace all the t parts by its equivalent u parts

OpenStudy (amistre64):

since we define u = t^2 + 7 ; wherever you see a t^2 + 7, replace it with a "u" dt has to be replaced by du .... but an equivalent form of it since u = t^2 + 7 , we can derive it and solve for dt du/dt = 2t du = 2t dt dt/2t = du is good enough

OpenStudy (anonymous):

When do we plug the t parts back in?

OpenStudy (amistre64):

replace all that parts:\[\int t(t^2+7)^2dt\to \ \int t(u)^2\frac{du}{2t}\]

OpenStudy (amistre64):

after we integrate up the u fillins, we replace the u parts by their t parts

OpenStudy (anonymous):

Integrate the second expression you put? Correct?

OpenStudy (amistre64):

itegrate either of them, but yes, the second one is spose to be the easier looking one

OpenStudy (anonymous):

\[\int\limits t^5 +14t^3 + 49t\ dt=\frac {t^6}{6}+14\frac{t^4}{4}+49\frac{t^2}{2}+C\]

OpenStudy (amistre64):

we should end up with:\[\frac{1}{6}u^3\] and replaceing the u parts; we know u = t^2 + 7\[\frac{1}{6}(t^2+7)^3\]

OpenStudy (anonymous):

ya thats what I got thanks!

OpenStudy (amistre64):

dont forget the +C arbitrary constant at the end

OpenStudy (anonymous):

\[\frac16 (t^2+7)^3=\frac {t^6}{ 6}+\frac{7t^4}{2}+\frac{49t^2}{2}+\frac{343}{6}+C\]

OpenStudy (anonymous):

\[\frac {343} 6+C=C_1\]